# Do Language Models Encode Collaborative Signals in Recommendation? Findings and Potentials

### v2

@@ Abstract
Recent studies empirically indicate that language models (LMs) encode rich world knowledge beyond mere semantics, attracting significant attention across various fields. However, in the recommendation domain, it remains uncertain whether LMs implicitly encode user preference information. Contrary to prevailing understanding that LMs and traditional recommenders learn two distinct representation spaces due to the huge gap in language and behavior modeling objectives, this work re-examines such understanding and explores extracting a recommendation space directly from the language representation space. Surprisingly, our findings demonstrate that item representations, when linearly mapped from advanced LM representations, yield superior recommendation performance. This outcome suggests the possible homomorphism between the advanced language representation space and an effective item representation space for recommendation, implying that collaborative signals may be implicitly encoded within LMs. Motivated by the finding of homomorphism, we explore the possibility of designing advanced collaborative filtering (CF) models purely based on language representations without ID-based embeddings. To be specific, we incorporate several crucial components (i.e., a multilayer perceptron (MLP), graph convolution, and contrastive learning (CL) loss function) to build a simple yet effective model, with the language representations of item textual metadata (i.e., title) as the input. Empirical results show that such a simple model can outperform leading ID-based CF models on multiple datasets, which sheds light on using language representations for better recommendation. Moreover, we systematically analyze this simple model and find several key features for using advanced language representations: a good initialization for item representations, superior zero-shot recommendation abilities in new datasets, and being aware of user intention. Our findings highlight the connection between language modeling and behavior modeling, which can inspire both natural language processing and recommender system communities  [^1] Codes are available at https://github.com/LehengTHU/AlphaRec.
@@ 1 Introduction
Currently, one prevailing understanding holds that general LMs and traditional recommenders (e.g., collaborative filtering models (Koren et al., 2009; He et al., 2021)) encode distinct representation spaces — one for language space and the other for behavior space — but they offer the potential to enhance each other in downstream recommendation tasks (Liao et al., 2024). Specifically, on the one hand, when using LMs as recommenders to directly output items of interest, aligning the language space with the behavior space can significantly improve the recommendation performance (Lin et al., 2023a; Vats et al., 2024; Xu et al., 2024). Various alignment strategies are proposed, including fine-tuning LMs with user behavior data (Zhang et al., 2023d; Bao et al., 2023; Geng et al., 2022; Cui et al., 2022; Lin et al., 2023b), incorporating embeddings from traditional recommenders as a new modality of LMs (Liao et al., 2024; Zhang et al., 2023e; Yang et al., 2023), and extending the vocabulary of LMs with item tokens (Zhu et al., 2023; Zheng et al., 2023; Rajput et al., 2023; Zhai et al., 2024). On the other hand, when using LMs as the enhancer to represent item content (e.g., text metadata), traditional recommenders greatly benefit from text embeddings (Yuan et al., 2022; 2023; Li et al., 2023c; Hou et al., 2024a; Liu et al., 2024a), semantic and reasoning information (Wei et al., 2024; Ren et al., 2024b; Xi et al., 2023), and generated user behaviors (Zhang et al., 2023b; c). Despite these efforts, explicit studies of the relationship between language and behavior spaces remain largely unexplored in the recommendation domain.
@@ 2.2 Item Representation Generation
LM-based generator. Beyond ID information, another research line (Pazzani & Billsus, 2007; Covington et al., 2016; Liu et al., 2024b; Zhang et al., 2024a; Liu et al., 2023b) explores using the text metadata of item $i$ (e.g., titles, descriptions) as pre-existing features $\mathbf{x}_{i}$. These features are fed into the LM-based generator, typically a combination of two subsequent components: (1) A frozen LM to extract $i$’s language representation $\mathbf{z}_{i}$, such as the encoder-only LMs like BERT-style models (Devlin et al., 2019; Liu et al., 2019), the decoder-only LMs like Llama-style autoregressive models (Touvron et al., 2023b; Jiang et al., 2023) and OpenAI text embedding models (Neelakantan et al., 2022); and (2) A trainable projector to map $\mathbf{z}_{i}$ into the final representation $\mathbf{e}_{i}$, often using layers like graph convolution layers (He et al., 2021). Although such LM-based generators have been explored to enrich the item representations in literatures (Yuan et al., 2023; Li et al., 2023c; Ren et al., 2024b), few studies have demonstrated that they can solely outperform ID-based generators in recommendation tasks. Worse still, the relationship between the LM-based language space and the behavior space remain largely unexplored in the recommendation domain.
@@ 4.2 Empirical Findings
Baselines. We compare AlphaRec with leading ID-based CF baselines, to assess the effectiveness of adopting advanced language representations. We do not consider baselines using LMs as recommenders for two practical reasons: the huge inference cost on datasets with millions of interactions and the task limitation of candidate selection (Liao et al., 2024) or next item prediction (Zheng et al., 2023). In addition to classic baselines introduced in section 3.2, we consider two categories of leading ID-based CF baselines, CL-based CF methods: SGL (Wu et al., 2021), BC Loss (Zhang et al., 2022), XSimGCL (Yu et al., 2024) and LM-enhanced CF methods: KAR (Xi et al., 2023), RLMRec (Ren et al., 2024b). See more details about baselines in Appendix C.2.
Recommendation capabilities (RQ3). Table 3 presents performance comparison. The best-performing methods are bold, while the second-best methods are underlined. We observe that:
@@
  • Advanced language representations shows strong potentials for recommendation, which can be unleashed by appropriate model design. AlphaRec consistently outperforms leading CF baselines by a large margin across all metrics on all datasets, with an improvement ranging from 6.79% to 9.75% on Recall@20 compared to the best baseline. Moreover, as shown in Figure 3(a), each component contributes positively (see more ablation results in Appendix C.3) in unleashing the power of language representations. Specifically, the performance degradation caused by replacing the MLP with a linear weight matrix (w/o MLP) indicates that nonlinear transformations can extract the implicit user preference similarities encoded in the language representation space more effectively. Besides, the performance also drops from replacing InfoNCE loss (Wu et al., 2022) with BPR loss (Rendle et al., 2012) (w/o CL) and removing the graph convolution (w/o GCN) suggests that explicitly modeling the collaborative relationships through the loss function and model architecture can further enhance recommendation performance. These findings suggest that the power of advanced language representations can be unleashed by carefully designing the model, showcasing the potential to surpass prevailing ID-based recommenders.
  • The incorporation of advanced language representations can benefit traditional ID-based CF methods. We note that two LM-enhanced CF methods, KAR and RLMRec, both show improvements over the most advanced CF methods. Nevertheless, the combination of ID-based embeddings and language representations in these methods does not yield higher results than purely language-representation-based AlphaRec. We attribute this phenomenon to their naive design for the combination ID-based embeddings and language representations, which is also highlighted by previous works (Yuan et al., 2023; Zhang et al., 2024c).
@@ 5.2 Zero-shot Ability (Potential 2)
Experimental settings. In zero-shot recommendation, there is no item or user overlap between the training set and test set (Ding et al., 2021; Zhang et al., 2024b), which is different from the research line of cross-domain recommendation (Zhu et al., 2021). We jointly train AlphaRec on three source datasets (i.e., Books, Movies & TV, and Video Games), while testing it on three completely new target datasets (i.e., Movielens-1M (Harper & Konstan, 2016), Book Crossing (Lee et al., 2019), and Amazon Industrial & Scientific (Ni et al., 2019)) without further training on these new datasets. (see more details about training on multiple datasets in Appendix D.2.1). Due to the lack of zero-shot recommenders in general CF, we slightly modify the two zero-shot methods in the sequential recommendation (Wang et al., 2019a), ZESRec (Hou et al., 2022) and UniSRec (Hou et al., 2022), as baselines. We also incorporate two strategy-based CF methods (i.e., Random and Pop) and one method using the large language model (LLM) as zero-shot recommender (i.e., LLMRank (Hou et al., 2024b)) (see more details about baselines in Appendix D.2.2).
@@ 5.3 Intention-aware Ability (Potential 3)
The language understanding ability in advanced language representations enables recommenders to perceive user intentions and refine recommendations. As shown in Table 5, the introduction of user intention (w Intention) significantly refines the recommendations of the pre-trained AlphaRec (w/o Intention). Moreover, AlphaRec outperforms the baseline model TEM (Bi et al., 2020) by a large margin, even without additional training on search tasks. We further conduct a case study on MovieLens-1M to demonstrate how AlphaRec captures the user intention (see more case study results in Appendix D.3.3). Additionally, the intention-aware ability also benefits from user historical interests. Figure 4(b) depicts the effect of $\alpha$. The $\alpha$ controls the strength of user intention, where $\alpha=0$ denotes that the user intention is neglected and $\alpha=1$ denotes that the user historical interest is ignored. The convex curve in Figure 4(b) suggests that both user historical interests and user intention play vital roles. The above findings suggest the potential of adopting advanced language representations to perceive text-based user intentions and refining recommendations (see more details in Appendix D.3).
@@ 6 Limitations
There are several limitations unaddressed in this paper. On the one hand, while we explore the relationship between language spaces and behavior spaces through empirical results, there is no theoretical guarantee. On the other hand, although we investigate the potential of advanced language representations for recommendation, we do not design any new components for CF models.
@@ 7 Conclusion
In this paper, we explored the relationship between language space and behavior spaces for recommendation, and explored the potential for using language representations for recommendation. Empirical results suggest the possible presence of homomorphism between advanced LMs representation spaces and an effective item representation space for recommendation. Inspired by this finding, we discussed how to unleash the power of advanced language representations by developing a simple yet effective CF model called AlphaRec. Moreover, by systematically analyzing AlphaRec, we explored the potentials of advanced language representations: fast convergence, zero-shot ability, and intention-aware ability. Possible future work will involve exploring the space relationship from both theoretical and multimodal perspectives. We believed that this paper sheds light on rethinking the connection between language modeling and user behavior modeling, benefiting both natural language processing and recommender system communities.
@@ Appendix A Related Works
Representations in LMs. The impressive capabilities demonstrated by LMs across various tasks raise a wide concern about what they have learned in the representation space. Linear methods (e.g., linear probing (Merullo et al., 2023) and linear mapping (Alain & Bengio, 2017)) are important and effective approaches for interpreting and analyzing representations of LMs (Ravichander et al., 2021). The main idea of linear methods is simple: training linear classifiers to predict some specific attributes or concepts (e.g., lexical structure (Vulic et al., 2020) ) from the representations in the hidden layers of LMs, or transforming language representations into another feature space with a linear matrix. A high result of linear methods (e.g., classification accuracy on the out-of-sample test set) tends to imply relevant information has been implicitly encoded in the representation space of LMs, although this does not imply LMs directly use these representations (Ravichander et al., 2021; Gurnee & Tegmark, 2023). Recent studies empirically demonstrate that concepts such as color (Patel & Pavlick, 2022), game states (Li et al., 2023a). and geographic position are encoded in LMs. Furthermore, these concepts may even be linearly encoded in the representation space of LMs (Li et al., 2023a; Park et al., 2023).
@@
Collaborative filtering. Collaborative filtering (CF) (Ren et al., 2024a) is an advanced technique in modern recommender systems. The prevailing CF methods tend to adopt an ID-based paradigm, where users and items are typically represented as one-hot vectors, with an embedding table used for lookup (Koren et al., 2009). Usually, these embedding parameters are learned by optimizing specific loss functions to reconstruct the history interaction pattern (Rendle et al., 2012). Recent advances in CF mainly benefit from two aspects, graph convolution (Wu et al., 2023b) and contrastive learning (Ren et al., 2024a). These CF models exhibit superior recommendation performance by conducting the embedding propagation (Wang et al., 2019b; He et al., 2021) and applying contrastive learning objectives (Wu et al., 2021; Cai et al., 2023; Yu et al., 2024). However, although effective, these methods are still limited, due to the ID-based paradigm. Since one-hot vectors contain no feature information beyond being identifiers, it is challenging to transfer pre-trained ID embeddings to other domains (Hou et al., 2022) or to leverage leading techniques from computer vision (CV) and natural language processing (NLP) (Yuan et al., 2023).
@@
LMs for recommendation. The remarkable language understanding and reasoning ability shown by LMs has attracted extensive attention in the field of recommendation. The application of LMs in recommendation can be categorized into three main approaches: LM-enhanced recommendation, LM as the modality encoder, and LM-as recommender. The first research direction, LM-enhanced recommendation, focuses on empowering traditional recommenders with the semantic representations from LMs (Xi et al., 2023; Ren et al., 2024b; Wei et al., 2024; Geng et al., 2024; Chen, 2023; Wang et al., 2024; Hou et al., 2023; Mao et al., 2023; Qiu et al., 2021; Zhang et al., 2024c). Specifically, these methods introduce representations from LMs as additional features for traditional ID-based recommenders, to capture complicated user preferences. The second research line lies in adopting the LM as the text modality encoder, which is also known as a kind of modality-based recommendation (MoRec) (Yuan et al., 2023; Li et al., 2023c). These methods tend to train the LM as the text modality encoder together with the traditional recommender. In previous studies, BERT-style LMs are widely used as the text modality encoder. The third research line fails in directly using LMs as the recommender and recommends items in a text generation paradigm. Early attempts focus on adopting in-context learning (ICL) (Dong et al., 2022) and prompting pre-trained LMs (Hou et al., 2024b; Liu et al., 2023a; Dai et al., 2023; Gao et al., 2023). However, such naive methods tend to yield poor performance compared to traditional models. Therefore, recent studies concentrate on fine-tuning LMs on recommendation-related corpus (Bao et al., 2023; Zhang et al., 2023d; Lin et al., 2023b; Cui et al., 2022; Liu et al., 2024b; Hua et al., 2023; Chen et al., 2024) and align the LMs with the representations from traditional recommenders as the additional modality (Liao et al., 2024; Zhang et al., 2023e; Yang et al., 2023; Li et al., 2023d; Kong et al., 2024).

### v3

# Language Representations Can be What Recommenders Need: Findings and Potentials

@@ Abstract
Recent studies empirically indicate that language models (LMs) encode rich world knowledge beyond mere semantics, attracting significant attention across various fields. However, in the recommendation domain, it remains uncertain whether LMs implicitly encode user preference information. Contrary to prevailing understanding that LMs and traditional recommenders learn two distinct representation spaces due to the huge gap in language and behavior modeling objectives, this work re-examines such understanding and explores extracting a recommendation space directly from the language representation space. Surprisingly, our findings demonstrate that item representations, when linearly mapped from advanced LM representations, yield superior recommendation performance. This outcome suggests the possible homomorphism between the advanced language representation space and an effective item representation space for recommendation, implying that collaborative signals may be implicitly encoded within LMs. Motivated by the finding of homomorphism, we explore the possibility of designing advanced collaborative filtering (CF) models purely based on language representations without ID-based embeddings. To be specific, we incorporate several crucial components (i.e., a multilayer perceptron (MLP), graph convolution, and contrastive learning (CL) loss function) to build a simple yet effective model, with the language representations of item textual metadata (i.e., title) as the input. Empirical results show that such a simple model can outperform leading ID-based CF models on multiple datasets, which sheds light on using language representations for better recommendation. Moreover, we systematically analyze this simple model and find several key features for using advanced language representations: a good initialization for item representations, superior zero-shot recommendation abilities in new datasets, and being aware of user intention. Our findings highlight the connection between language modeling and behavior modeling, which can inspire both natural language processing and recommender system communities.
@@ 1 Introduction
Currently, one prevailing understanding holds that general LMs and traditional recommenders (e.g., collaborative filtering models (Koren et al., 2009; He et al., 2021)) encode distinct representation spaces — one for language space and the other for behavior space — but they offer the potential to enhance each other in downstream recommendation tasks (Liao et al., 2024). Specifically, on the one hand, when using LMs as recommenders to directly output items of interest, aligning the language space with the behavior space can significantly improve the recommendation performance (Lin et al., 2023a; Vats et al., 2024; Xu et al., 2024). Various alignment strategies are proposed, including fine-tuning LMs with user behavior data (Zhang et al., 2023e; Bao et al., 2023; Geng et al., 2022; Cui et al., 2022; Lin et al., 2023b), incorporating embeddings from traditional recommenders as a new modality of LMs (Liao et al., 2024; Zhang et al., 2023f; Yang et al., 2023; Tennenholtz et al., 2024), and extending the vocabulary of LMs with item tokens (Zhu et al., 2023; Zheng et al., 2023; Rajput et al., 2023; Zhai et al., 2024). On the other hand, when using LMs as the enhancer to represent item content (e.g., text metadata), traditional recommenders greatly benefit from text embeddings (Yuan et al., 2022; 2023; Li et al., 2023c; Hou et al., 2024a; Liu et al., 2024b), semantic and reasoning information (Wei et al., 2024; Ren et al., 2024b; Xi et al., 2023), and generated user behaviors (Zhang et al., 2023b; d). Despite these efforts, explicit studies of the relationship between language and behavior spaces remain largely unexplored in the recommendation domain.
@@ 2.2 Item Representation Generation
LM-based generator. Beyond ID information, another research line (Pazzani & Billsus, 2007; Covington et al., 2016; Liu et al., 2024c; Zhang et al., 2024b; Liu et al., 2023b) explores using the text metadata of item $i$ (e.g., titles, descriptions) as pre-existing features $\mathbf{x}_{i}$. These features are fed into the LM-based generator, typically a combination of two subsequent components: (1) A frozen LM to extract $i$’s language representation $\mathbf{z}_{i}$, such as the encoder-only LMs like BERT-style models (Devlin et al., 2019; Liu et al., 2019), the decoder-only LMs like Llama-style autoregressive models (Touvron et al., 2023b; Jiang et al., 2023) and OpenAI text embedding models (Neelakantan et al., 2022); and (2) A trainable projector to map $\mathbf{z}_{i}$ into the final representation $\mathbf{e}_{i}$, often using layers like graph convolution layers (He et al., 2021). Although such LM-based generators have been explored to enrich the item representations in literatures (Yuan et al., 2023; Li et al., 2023c; Ren et al., 2024b), few studies have demonstrated that they can solely outperform ID-based generators in recommendation tasks. Worse still, the relationship between the LM-based language space and the behavior space remain largely unexplored in the recommendation domain.
@@ 4.2 Empirical Findings
Baselines. We compare AlphaRec with leading ID-based CF baselines, to assess the effectiveness of adopting advanced language representations. In addition to classic baselines introduced in section 3.2, we consider two categories of leading ID-based CF baselines, CL-based CF methods: SGL (Wu et al., 2021), BC Loss (Zhang et al., 2022), XSimGCL (Yu et al., 2024) and LM-enhanced CF methods: KAR (Xi et al., 2023), RLMRec (Ren et al., 2024b). We also consider a variant of XSimGCL (i.e., XSimGCLt) where ID embeddings are replaced with language representations, and directly using language representations $\mathbf{z}_{u}$ and $\mathbf{z}_{i}$ for recommendation without training (i.e., EMB-KNN). See baseline details in Appendix C.2.
Recommendation capabilities (RQ3). Table 3 presents performance comparison. To alleviate the possible information leakage problem that pretraining data of LMs may include the training data, we also conduct experiments on a newly published dataset Amazon 2023 in Appendix C.6. The best-performing methods are bold, while the second-best methods are underlined. We observe that:
@@
  • Advanced language representations shows strong potentials for recommendation, which can be unleashed by appropriate model design. AlphaRec consistently outperforms leading CF baselines by a large margin across all metrics on all datasets, with an improvement ranging from 6.79% to 9.75% on Recall@20 compared to the best baseline. Moreover, as shown in Figure 3(a), each component contributes positively (see more ablation results and details in Appendix C.3) in unleashing the power of language representations. Moreover, simply replacing the ID embeddings in XSimGCL with language representations (i.e., XSimGCLt) does not bring a stable and remarkable improvement, which is consistent with previous findings (Yuan et al., 2023; Li et al., 2023c). This phenomenon highlights the importance of appropriate model design beyond simple input feature replacement. Above findings suggest that the power of advanced language representations can be unleashed by carefully designing the model, showcasing the potential to surpass prevailing ID-based recommenders.
  • The incorporation of advanced language representations can benefit traditional ID-based CF methods. We note that two LM-enhanced CF methods, KAR and RLMRec, both show improvements over the most advanced CF methods. Nevertheless, the combination of ID-based embeddings and language representations in these methods does not yield higher results than purely language-representation-based AlphaRec. We attribute this phenomenon to their naive design for the combination ID-based embeddings and language representations, which is also highlighted by previous works (Yuan et al., 2023; Zhang et al., 2024d).
@@ 5.2 Zero-shot Ability (Potential 2)
Experimental settings. In zero-shot recommendation, there is no item or user overlap between the training set and test set (Ding et al., 2021; Zhang et al., 2024c), which is different from the research line of cross-domain recommendation (Zhu et al., 2021). We jointly train AlphaRec on three source datasets (i.e., Books, Movies & TV, and Video Games), while testing it on three completely new target datasets (i.e., Movielens-1M (Harper & Konstan, 2016), Book Crossing (Lee et al., 2019), and Amazon Industrial & Scientific (Ni et al., 2019)) without further training on these new datasets. (see more details about training on multiple datasets in Appendix D.2.1). Due to the lack of zero-shot recommenders in general CF, we slightly modify three zero-shot methods in the sequential recommendation (Wang et al., 2019a), ZESRec (Hou et al., 2022), UniSRec (Hou et al., 2022), and VQ-Rec (Hou et al., 2023), as baselines. We also incorporate two strategy-based CF methods (i.e., Random and Pop) and one method using the large language model (LLM) as zero-shot recommender (i.e., LLMRank (Hou et al., 2024b)) (see more details about baselines in Appendix D.2.2).
@@ 5.3 Intention-aware Ability (Potential 3)
The language understanding ability in advanced language representations enables recommenders to perceive user intentions and refine recommendations. As shown in Table 5, the introduction of user intention (w Intention) significantly refines the recommendations of the pre-trained AlphaRec (w/o Intention). Moreover, AlphaRec outperforms the baseline model TEM (Bi et al., 2020), even without additional training on search tasks. We further conduct a case study on MovieLens-1M to demonstrate how AlphaRec captures the user intention (see more examples in Appendix D.3.3). Additionally, the intention-aware ability benefits from user historical interests. Figure 4(b) depicts the effect of $\alpha$, which controls the strength of user intention. Here $\alpha=0$ denotes that the user intention is neglected and $\alpha=1$ denotes that the user historical interest is ignored. The convex curve in Figure 4(b) suggests that both user interests and user intention play vital roles. Above findings highlight the potential of adopting advanced language representations to perceive text-based user intentions and refine recommendations (see more details in Appendix D.3).
@@ 6 Limitations
There are several limitations unaddressed in this paper. First, this research lacks theoretical guarantee, and we do not contribute to designing any new components for CF models. Second, AlphaRec lacks a personalized design, since there is only a single MLP introduced for all users. The fixed user intention in Section 5.3 is less practical, lacking exploration on personalized user intentions.
@@ 7 Conclusion
In this paper, we explored the relationship between language space and behavior spaces for recommendation, and explored the potential for using language representations for recommendation. Empirical results suggest the possible presence of homomorphism between advanced LMs representation spaces and an effective item representation space for recommendation. Inspired by this finding, we discussed how to unleash the power of advanced language representations by developing a simple yet effective CF model called AlphaRec. Moreover, by systematically analyzing AlphaRec, we explored the potentials of advanced language representations: good initialization for item representations, zero-shot ability, and intention-aware ability. Possible future work will involve exploring the space relationship from both theoretical and multimodal perspectives. We believed that this paper sheds light on rethinking the connection between language modeling and user behavior modeling, benefiting both natural language processing and recommender system communities.
Acknowledgments
This research is supported by the National Science and Technology Major Project (2023ZD0121102), the NExT Research Center, National Natural Science Foundation of China (92270114).
@@ Appendix A Related Works
Representations in LMs. The impressive capabilities demonstrated by LMs across various tasks raise a wide concern about what they have learned in the representation space. Linear methods (e.g., linear probing (Merullo et al., 2023) and linear mapping (Alain & Bengio, 2017)) are important and effective approaches for interpreting and analyzing representations of LMs (Ravichander et al., 2021). The main idea of linear methods is simple: training linear classifiers to predict some specific attributes or concepts (e.g., lexical structure (Vulic et al., 2020) ) from the representations in the hidden layers of LMs, or transforming language representations into another feature space with a linear matrix. A high result of linear methods (e.g., classification accuracy on the out-of-sample test set) tends to imply relevant information has been implicitly encoded in the representation space of LMs, although this does not imply LMs directly use these representations (Ravichander et al., 2021; Gurnee & Tegmark, 2023; Fang et al., 2025). Recent studies empirically demonstrate that concepts such as color (Patel & Pavlick, 2022), game states (Li et al., 2023a). and geographic position are encoded in LMs. Furthermore, these concepts may even be linearly encoded in the representation space of LMs (Li et al., 2023a; Park et al., 2023).
@@
Collaborative filtering. Collaborative filtering (CF) (Ren et al., 2024a) is an advanced technique in modern recommender systems. The prevailing CF methods tend to adopt an ID-based paradigm, where users and items are typically represented as one-hot vectors, with an embedding table used for lookup (Koren et al., 2009). Usually, these embedding parameters are learned by optimizing specific loss functions to reconstruct the history interaction pattern (Rendle et al., 2012). Recent advances in CF mainly benefit from two aspects, graph convolution (Wu et al., 2023b) and contrastive learning (Ren et al., 2024a; Zhang et al., 2023c). These CF models exhibit superior recommendation performance by conducting the embedding propagation (Wang et al., 2019b; He et al., 2021) and applying contrastive learning objectives (Wu et al., 2021; Cai et al., 2023; Yu et al., 2024; Zhang et al., 2024a). However, although effective, these methods are still limited, due to the ID-based paradigm. Since one-hot vectors contain no feature information beyond being identifiers, it is challenging to transfer pre-trained ID embeddings to other domains (Hou et al., 2022) or to leverage leading techniques from computer vision (CV) and natural language processing (NLP) (Yuan et al., 2023).
@@
LMs for recommendation. The remarkable language understanding and reasoning ability shown by LMs has attracted extensive attention in the field of recommendation. The application of LMs in recommendation can be categorized into three main approaches: LM-enhanced recommendation, LM as the modality encoder, and LM-as recommender. The first research direction, LM-enhanced recommendation, focuses on empowering traditional recommenders with the semantic representations from LMs (Xi et al., 2023; Ren et al., 2024b; Wei et al., 2024; Geng et al., 2024; Chen, 2023; Wang et al., 2024; Hou et al., 2023; Mao et al., 2023; Qiu et al., 2021; Zhang et al., 2024d; Liu et al., 2024a; d). Specifically, these methods introduce representations from LMs as additional features for traditional ID-based recommenders, to capture complicated user preferences. The second research line lies in adopting the LM as the text modality encoder, which is also known as a kind of modality-based recommendation (MoRec) (Yuan et al., 2023; Li et al., 2023c). These methods tend to train the LM as the text modality encoder together with the traditional recommender. In previous studies, BERT-style LMs are widely used as the text modality encoder. The third research line fails in directly using LMs as the recommender and recommends items in a text generation paradigm. Early attempts focus on adopting in-context learning (ICL) (Dong et al., 2022) and prompting pre-trained LMs (Hou et al., 2024b; Liu et al., 2023a; Dai et al., 2023; Gao et al., 2023). However, such naive methods tend to yield poor performance compared to traditional models. Therefore, recent studies concentrate on fine-tuning LMs on recommendation-related corpus (Bao et al., 2023; Zhang et al., 2023e; Lin et al., 2023b; Cui et al., 2022; Liu et al., 2024c; Hua et al., 2023; Chen et al., 2024; Ding et al., 2024) and align the LMs with the representations from traditional recommenders as the additional modality (Liao et al., 2024; Zhang et al., 2023f; Yang et al., 2023; Li et al., 2023d; Kong et al., 2024).